Please Help!!
A pharmicist wishes to mix a solution that is 5% minoxisil. She has on hand 100 ml of a 3% solution & wishes to add some 7% solution to obtain the desired 5% solution. How much 7% solution should she add??

Answer by scott8148(6628)

let x="ml of added 7% solution"

(3%)(100) + (7%)(x) = (5%)(100 + x)

multiplying by 100 to clear percentages ___ 300 + 7x = 500 + 5x

2x = 200 ___ x = 100


this makes sense ...
___ you would expect that mixing equal volumes of different concentrations would give the average of the concentrations